OpenStudy (anonymous):

Radium -226 has a half life of about 110 years. After 350 years how much radium would be left if you started with 1700 mg?

OpenStudy (anonymous):

\[C _{f}=C _{o}\left(\begin{matrix}\frac{ 1 }{ 2 }\end{matrix}\right)^{\frac{ time }{ half-life }}\]

OpenStudy (anonymous):

Another formula for continuous decay is \[C _{f}=C _{o}e ^{-\lambda*t}\] Where \[\lambda=\frac{ \ln2 }{ half-life }\]

Directrix (directrix):

y = 1700 * (.5)^(350/110) y = 1700 * .1102 y = ? @Tyndall25 Post what you get, okay? Also, check out the online half-life calculator at this link: http://www.calculator.net/half-life-calculator.html

OpenStudy (anonymous):

187.34

Directrix (directrix):

That's what I got.